OpenStudy (anonymous):

r, s, and t are roots of Ax^3+Bx^2+Cx+D=0. Express 1/r+1/s+1/t in terms of A, B, C, and D.

OpenStudy (anonymous):

Viete's equation.

OpenStudy (anonymous):

i am going to make a guess that it is \[\frac{\frac{c}{a}}{-\frac{d}{a}}=-\frac{c}{d}\] lets see if it is right

OpenStudy (anonymous):

\[\frac{1}{r}+\frac{1}{s}+\frac{1}{t}=\frac{rs+rt+st}{rst}\] looks good

OpenStudy (anonymous):

since we have\(rst=-\frac{d}{a}\) and \(rs+rt+st=\frac{c}{a}\) by those vieta's formulae

OpenStudy (anonymous):

Oh, you substitute the values. I see now. Thank you!

OpenStudy (anonymous):

yw

OpenStudy (aravindg):

Hi, I can see that you are a new member here at OpenStudy so, I would first like to say Welcome to OpenStudy, I would like to point you to the chat pods these chat pods are where you can make new friends and talk to new people just like you, I would also like to emphasize our "NON-CHEATING POLICY" we ask here at OpenStudy that you do NOT post exam/Test questions if you are caught doing this we will notify your current school, I would also like to ask you to check out the OpenStudy "Code of Conduct" http://openstudy.com/code-of-conduct Please read this carefully and thoroughly. Welcome to the OpenStudy Community! Ambassador